# Nested, asymmetric H–He circumstellar shells in the Type Icn/Ibn SN 2024abvb

### v1

@@ 3.4 Spectral Evolution
There is clear evidence of spectral evolution in SN 2024abvb spectra. During the early phases (up to $\phi=5.7$), the spectra are dominated by prominent He i, C ii and O ii features. Between $\phi=-2.0$ and $\phi=5.7$, these lines exhibit relatively slow evolution, with their velocities remaining nearly constant over time. The minimum velocities measured for different ions are approximately $\sim 130$ km s-1 for O ii, $\sim 100$ km s-1 for C ii and around 120 km s-1 for He i. By $\phi=17.2$, the C ii and O ii features have disappeared, indicating significant changes in the ionisation conditions and/or chemical composition of the CSM. However, several He i lines, such as $\lambda 5015$, $\lambda 5876$, $\lambda 6678$ and $\lambda 7065$ remain detectable, albeit with reduced strength.
Figure 7: Same as Fig.6 but for O ii.
Figures 6, 7 and 8 show the temporal evolution of these lines at $\phi=5.7$ and $\phi=17.2$. Despite the overall similarity observed in these P-Cygni profiles at 5.7d, not all emission peaks are centred at the rest-frame wavelength. In several cases, the emission components exhibit redshifted peak offsets of $\sim 50-100$ km s-1, which could be attributed to asymmetries in the density distribution of ejecta.
Figure 8: Same as Fig.6 but for C ii.
Simultaneously, a decline in the strength of Fe ii features, particularly multiplet 42 by 17.2d, indicates a rapid evolution of the circumstellar environment and increasing optical depth in the NUV due to the emergence of iron-rich layers. In contrast, the Ca ii NIR triplet lines are visible at this epoch, displaying expansion velocities of $\sim 100$ km s-1, roughly consistent with the velocities of He i, C ii and O ii lines.
Although the XShooter spectra at both epochs are primarily dominated by narrow features (aside from broad bases to some lines, which we attribute to electron scattering), there is also evidence of some higher velocity material. In particular, the OI $\lambda$7774 line at 17.2d (Fig. 9) shows a distinct absorption component with a blue edge extending up to 2000 km s-1. This is also observed for the Ca ii NIR triplet. This broad absorption appears separate from the narrower, low-velocity component and suggests the presence of an outer shell of faster-moving ejecta. In the Sobolev approximation, the shape of these lines would suggest a density profile that is at least $\rho(r)\propto r^{-2}$ or steeper, as well as a relatively low value for $\tau_{0}$, the optical depth at the photosphere. The latter is also consistent with the absence of strong emission components to these lines. This would disfavour a dense shell origin for the CSM surrounding SN 2024abvb.
@@ 3.5 Comparison to Ibn/Icn
In Fig. 11 we compare two XShooter epochs of SN 2024abvb with the prototypical Type Icn SN 2019hgp (Gal-Yam et al., 2022), the transitional Type Ibn SN 2010al observed with high-resolution spectroscopy (Pastorello et al., 2015c), and PS1–12sk, a Type Ibn event that exploded at a large projected distance from its host galaxy (Sanders et al., 2013). The spectra of SN 2024abvb show less prominent O ii features than SN 2019hgp and no highly ionised carbon lines (e.g. C iii), particularly at wavelengths bluer than 5000 Å and usually observed in Icn events (Pellegrino et al., 2022). The C ii line velocities in SN 2024abvb are comparable to those measured in SN 2019hgp, although they are systematically slightly lower at similar evolutionary phases. A similar behaviour is observed when comparing He i velocities with those of SN 2010al, which lacks prominent O i and C ii features, or shows equivalent widths less than half of those measured in SN 2024abvb at both comparison epochs.
The helium line profiles of SN 2024abvb also differ from those typically observed in Type Ibn SNe, including PS1-12sk, as well as from the transitional case SN 2010al. In particular, SN 2024abvb lacks the broad emission component commonly seen in Ibn events, while the velocity of the narrow P-Cygni absorption is approximately a factor of two-three lower than that observed in typical Type Ibn SNe. Moreover, the spectra of SN 2024abvb exhibit a redder continuum than both Type Ibn and Icn events (see also Aster et al. 2026, submitted) and do not show the Fe ii pseudo-continuum commonly observed, albeit with varying strength, in Type Icn SNe (e.g. SN 2019hgp) and Type Ibn SNe (e.g. PS1-12sk). Taken together, these similarities and differences in ionic species, line strengths, and velocities support, from a phenomenological perspective, the classification of SN 2024abvb as a transitional Type Icn/Ibn event.
@@ 4.1 Interstellar polarisation
Determining the amount of interstellar polarisation (ISP) caused by dust grains in the line-of-sight is crucial to investigating the intrinsic SN polarisation. As Na ID absorption at host redshift is not present in the UVES or X-Shooter spectra, the extinction must be very low, as expected from the remote location of SN 2024abvb from its host galaxy (see Section 2.2. As such, the host contribution to the ISP is insignificant; it should be dominated by the Milky Way dust column.
With our VLT/FORS2 $BVRI$ polarimetry, we can use field stars to estimate the wavelength-dependent ISP. The field is very sparse, with only $\lesssim 5$ stars per band that are $\mathrm{S/N}>200$ needed for an accurate ISP estimate. However, most of them are found within 100 px from the edges of the image, and they appear to have higher polarisation than the ones nearer to the centre of the camera. The known instrumental polarisation increases sharply towards the edges (González-Gaitán et al., 2020), and while we correct for the effect, it is possible that the correction is not perfect at such locations. As such, we exclude these stars from measuring the ISP. Were they included the difference would be small as $B$ and $V$ would remain the same, and for $R$ and $I$ the ISP would increase by $P\lesssim 0.3\%$. We verified that the selected stars are suitable for measuring Galactic ISP. Following Tran (1995), stars must lie $\gtrsim 150$ pc above the Galactic plane to reliably probe the dust column, and astrometric solutions from Gaia Data Release 3 (Collaboration et al., 2022) show that the selected stars are well above this threshold.
@@ 4.2 Polarimetric evolution
The polarimetric evolution in $BVRI$, based on the VLT/FORS2 spectropolarimetry and broadband polarimetry as well as the NOT/ALFOSC $V$-band polarimetry, is shown in Fig. 13. The SN shows a consistent $P\sim 1\%$ near-peak in both FORS2 and ALFOSC data, but the polarisation decreases to $\lesssim 0.5$% at a week post-peak. After this, the polarisation increases again at a rotation of $50\deg$. At +22+22 d the polarisation degree is $\sim 1.5$% in $BV$, and at +27+27 d the value appears to have increased drastically to $P\sim 4$% in $BV$ while $RI$ exhibit more moderate values at $P\sim 2$%. The uncertainties of the last epoch are high, but it is clear that they have increased to a high degree at a physical rotation of $50\deg$ compared to near-peak. Further, the fact that both $BV$ are significantly higher than $RI$ indicates wavelength dependency. The observation was taken in prime conditions during 1% Lunar illumination, and the trend is not caused by significant scattering of Lunar photons in the atmosphere (see e.g. Pursiainen et al., 2023). Given the remote location, the background is also easy to model, and different methods for estimating it do not affect the results.
@@ 5 Discussion
To provide quantitative context for the discussion below, we begin by deriving order-of-magnitude estimates of the characteristic CSM radii and masses implied by the spectroscopic and polarimetric data. Assuming an ionised CSM with electron densities of $n_{e}\sim 10^{9}$$10^{10}$ cm-3 and characteristic radii $R\sim 10^{14}$$10^{15}$ cm, we can estimate the CSM mass at the order-of-magnitude level. Adopting a shell (or toroidal) geometry with a fractional thickness $\Delta R\sim 0.1$$0.3$ R (Smith, 2014; Moriya et al., 2014; Dessart et al., 2016) and accounting for a limited solid-angle covering factor, we infer CSM masses of $M_{\rm CSM}\sim 10^{-2}$ M per individual structure. The presence of multiple kinematically distinct components implies a total CSM mass of $M_{\rm CSM,total}\sim 10^{-1}$ M, although this value remains highly uncertain and model dependent.
@@ 5.1 SN 2024abvb as a thermonuclear supernova
In Fig. 14, we compare the optical lightcurve of SN 2024abvb to a number of Type Ia templates (Nugent et al., 2002), shifted so that the maxima of their lightcurves coincide. It is clear that a normal Type Ia SN could not be hidden under the light curve of SN 2024abvb, as from around 20 days onward, we would see a difference in the light curve. Moreover, the SN would dominate the emission at the epoch of our second XShooter spectrum, and so we would expect to see underlying photospheric features from the SN. From a luminosity point of view, accommodating a sub-luminous SN 1991bg-like SN is more plausible; in such a scenario, the underlying SN would be significantly fainter than our observed light curve for all epochs where we have spectra, and approximately consistent with our upper limits on the SN magnitude at early and late times. Nevertheless, the presence of H in the CSM would prove challenging to explain. Similarly, a SN Iax could also be hidden below the CSM interaction as they are several magnitudes below both the detections and upper limits of SN 2024abvb at all phases (e.g. Foley et al., 2013; Srivastav et al., 2020). However, they are not found in remote locations and show an association to young stellar populations (Lyman et al., 2018). A similar argument applies to a .Ia SN, whose intrinsic light curve would be hidden beneath the luminosity produced by CSM interaction (e.g. Inserra et al., 2015; Dong et al., 2022; Padilla Gonzalez et al., 2024). Such SNe have been found in remote locations and are thought to be triggered by a detonation of a helium layer on a low-mass CO white dwarf, which would also accommodate the presence of He in the CSM.
@@ 5.3 SN 2024abvb from a massive star or binary system
Finally, the relative asymmetries we retrieve can be reconciled with a time-dependent mass-loss axis. In a binary/common-envelope context, the orientation and degree of asymmetry of successive ejection episodes can change because of orbital evolution, episodic disk formation and dissipation, or torque-induced precession of an accretion disk or outflow-launching structure (e.g. Ivanova et al., 2013; Tauris et al., 2015). Such mechanisms naturally produce concentric but non-coaxial tori whose apparent flattening and inclination vary with epoch; the outermost shells then record an earlier geometry, while inner shells trace later, possibly reoriented mass-loss. We therefore consider a physically conservative interpretation in which (a) repeated, binary-driven ejection episodes produce concentric toroidal shells, (b) transient colliding-wind shocks enable dust condensation in some episodes (providing the wavelength-dependent polarisation), and (c) changes in the instantaneous mass-loss symmetry axis (e.g. driven by precession or secular orbital reorientation) produce the observed sequence of inclinations and asymmetries. While WR binaries provide a direct observational analogue for nested dusty shells (Lau et al., 2022), detailed hydrodynamic modelling targeted at CE/post-CE mass loss in stripped progenitors is required to quantify whether the exact sequence of asymmetries and kinematics seen here can be reproduced.
@@ 6 Conclusions
We propose the following conservative interpretation. The data are consistent with episodic, binary-driven mass loss from a common envelope scenario leading to the explosion of an electron-capture supernova surrounded by multiple, concentric (ring-like) CSM tori with differing orientations. In this picture, the outer (earlier) tori are more oblate, while intermediate shells are more spherical and the innermost shells are again more aspherical, a sequence that can be produced by changes in the instantaneous mass-loss axis driven by orbital evolution, transient accretion disks, or torque-induced precession during common-envelope / post-common-envelope phases. Dust condensed in earlier, more distant tori can naturally explain the late-time, wavelength-dependent increase in polarisation (blue-weighted) when different dusty shells come into the line of sight or are illuminated by the fading transient. This interpretation is supported by empirical examples of nested dust/ring structures in massive colliding-wind systems and the qualitative consistency of our measured kinematics and polarimetric evolution with episodic tilted ejection. While, to our knowledge, this represents the first SN with spectroscopic evidence consistent with nested, concentric tori, detailed hydrodynamic and radiative-transfer models targeted at CE/post-CE mass loss in stripped progenitors are required to confirm whether the precise sequence of inclinations, asymmetries and kinematics can be reproduced.
@@ Appendix A Additional Tables and Figures
Figure 15: The Serkowski fit to the $BVRI$ ISP estimated based on the FORS2 broadband polarimetry. The best-fitting model (thick solid line), as well as model fits to randomised 1000 Monte Carlo resamplings, are shown.

### v2

@@ 3.4 Spectral Evolution
There is clear evidence of spectral evolution in SN 2024abvb spectra. During the early phases (up to $\phi=5.7$), the spectra are dominated by prominent He i, C $i$ and O $i$ features. Between $\phi=-2.0$ and $\phi=5.7$, these lines exhibit relatively slow evolution, with their velocities remaining nearly constant over time. The minimum velocities measured for different ions are approximately $\sim 130$ km s-1 for O $i$, $\sim 100$ km s-1 for C $i$ and around 120 km s-1 for He i. By $\phi=17.2$, the C $i$ and O $i$ features have disappeared, indicating significant changes in the ionisation conditions and/or chemical composition of the CSM. However, several He i lines, such as $\lambda 5015$, $\lambda 5876$, $\lambda 6678$ and $\lambda 7065$ remain detectable, albeit with reduced strength.
Figures 6, 15 and 16 show the temporal evolution of these lines at $\phi=5.7$ and $\phi=17.2$, while 17 also includes the last NIR epoch at $\phi=55.7$. Despite the overall similarity observed in these P-Cygni profiles at 5.7d, not all emission peaks are centred at the rest-frame wavelength. In several cases, the emission components exhibit redshifted peak offsets of $\sim 50-100$ km s-1, which could be attributed to asymmetries in the density distribution of ejecta.
Simultaneously, a decline in the strength of Fe $i$ features, particularly multiplet 42 by 17.2d, indicates a rapid evolution of the circumstellar environment and increasing optical depth in the NUV due to the emergence of iron-rich layers. In contrast, the Ca $i$ NIR triplet lines are visible at this epoch, displaying expansion velocities of $\sim 100$ km s-1, roughly consistent with the velocities of He i, C $i$ and O $i$ lines.
Although the XShooter spectra at both epochs are primarily dominated by narrow features (aside from broad bases to some lines, which we attribute to electron scattering), there is also evidence of some higher velocity material. In particular, the OI $\lambda$7774 line at 17.2d (Fig. 7) shows a distinct absorption component with a blue edge extending up to 2000 km s-1. This is also observed for the Ca $i$ NIR triplet. This broad absorption appears separate from the narrower, low-velocity component and suggests the presence of an outer shell of faster-moving ejecta. In the Sobolev approximation, the shape of these lines would suggest a density profile that is at least $\rho(r)\propto r^{-2}$ or steeper, as well as a relatively low value for $\tau_{0}$, the optical depth at the photosphere. The latter is also consistent with the absence of strong emission components to these lines. This would disfavour a dense shell origin for the CSM surrounding SN 2024abvb.
@@ 3.5 Comparison to Ibn/Icn
In Fig. 9 we compare two XShooter epochs of SN 2024abvb with the prototypical Type Icn SN 2019hgp (Gal-Yam et al., 2022), the transitional Type Ibn SN 2010al observed with high-resolution spectroscopy (Pastorello et al., 2015), and PS1–12sk, a Type Ibn event that exploded at a large projected distance from its host galaxy (Sanders et al., 2013). The spectra of SN 2024abvb show less prominent O $i$ features than SN 2019hgp and no highly ionised carbon lines (e.g. C iii), particularly at wavelengths bluer than 5000 Å and usually observed in Icn events (Pellegrino et al., 2022). The C $i$ line velocities in SN 2024abvb are comparable to those measured in SN 2019hgp, although they are systematically slightly lower at similar evolutionary phases. A similar behaviour is observed when comparing He i velocities with those of SN 2010al, which lacks prominent O i and C $i$ features, or shows equivalent widths less than half of those measured in SN 2024abvb at both comparison epochs.
The helium line profiles of SN 2024abvb also differ from those typically observed in Type Ibn SNe, including PS1-12sk, as well as from the transitional case SN 2010al. In particular, SN 2024abvb lacks the broad emission component commonly seen in Ibn events, while the velocity of the narrow P-Cygni absorption is approximately a factor of two-three lower than that observed in typical Type Ibn SNe. Moreover, the spectra of SN 2024abvb exhibit a redder continuum than both Type Ibn and Icn events (see Aster et al., 2026) and do not show the Fe $i$ pseudo-continuum commonly observed, albeit with varying strength, in Type Icn SNe (e.g. SN 2019hgp) and Type Ibn SNe (e.g. PS1-12sk). Taken together, these similarities and differences in ionic species, line strengths, and velocities support, from a phenomenological perspective, the classification of SN 2024abvb as a transitional Type Icn/Ibn event.
@@ 4.1 Interstellar polarisation
Determining the amount of interstellar polarisation (ISP) caused by dust grains in the line-of-sight is crucial to investigating the intrinsic SN polarisation. As Na ID absorption at host redshift is not present in the UVES or X-Shooter spectra, the extinction must be very low, as expected from the remote location of SN 2024abvb from its host galaxy (see Section $2.$ As such, the host contribution to the ISP is insignificant; it should be dominated by the Milky Way dust column.
With our VLT/FORS2 BVRIBVRI polarimetry, we can use field stars to estimate the wavelength-dependent ISP. The field is very sparse, with only $\lesssim 5$ stars per band that are $\mathrm{S/N}>200$ needed for an accurate ISP estimate. However, most of them are found within 100 px from the edges of the image, and they appear to have higher polarisation than the ones nearer to the centre of the camera. The known instrumental polarisation increases sharply towards the edges (González-Gaitán et al., 2020), and while we correct for the effect, it is possible that the correction is not perfect at such locations. As such, we exclude these stars from measuring the ISP. Were they included the difference would be small as $B$ and $V$ would remain the same, and for $R$ and II the ISP would increase by $P\lesssim 0.3\%$. We verified that the selected stars are suitable for measuring Galactic ISP. Following Tran (1995), stars must lie $\gtrsim 150$ pc above the Galactic plane to reliably probe the dust column, and astrometric solutions from Gaia Data Release 3 (Collaboration et al., 2022) show that the selected stars are well above this threshold.
@@ 4.2 Polarimetric evolution
The polarimetric evolution in BVRIBVRI, based on the VLT/FORS2 spectropolarimetry and broadband polarimetry as well as the NOT/ALFOSC $V$-band polarimetry, is shown in Fig. 11. The SN shows a consistent $P\sim 1\%$ near-peak in both FORS2 and ALFOSC data, but the polarisation decreases to $\lesssim 0.5$% at a week post-peak. After this, the polarisation increases again at a rotation of $50\deg$. At +22+22 d the polarisation degree is $\sim 1.5$% in BVBV, and at +27+27 d the value appears to have increased drastically to $P\sim 4$% in BVBV while RIRI exhibit more moderate values at $P\sim 2$%. The uncertainties of the last epoch are high, but it is clear that they have increased to a high degree at a physical rotation of $50\deg$ compared to near-peak. Further, the fact that both BVBV are significantly higher than RIRI indicates wavelength dependency. The observation was taken in prime conditions during 1% Lunar illumination, and the trend is not caused by significant scattering of Lunar photons in the atmosphere (see e.g. Pursiainen et al., 2023). Given the remote location, the background is also easy to model, and different methods for estimating it do not affect the results.
@@ 5 Discussion
To provide quantitative context for the discussion below, we begin by deriving order-of-magnitude estimates of the characteristic CSM radii and masses implied by the spectroscopic and polarimetric data. Assuming an ionised CSM with electron densities of $n_{e}\sim 10^{9}$$10^{10}$ cm-3 and characteristic radii $R\sim 10^{14}$$10^{15}$ cm, we can estimate the CSM mass at the order-of-magnitude level. Adopting a shell (or toroidal) geometry with a fractional thickness $\Delta R\sim 0.1$$0.3$ R (Smith, 2014; Moriya et al., 2014; Dessart et al., 2016) and accounting for a limited solid-angle covering factor, we infer CSM masses of $M_{\rm CSM}\sim 10^{-2}$ M per individual structure. The presence of multiple kinematically distinct components implies a total CSM mass of $M_{\rm CSM,total}\sim 10^{-1}$ M, although this value remains highly uncertain and model dependent. Such a value is comparable to the CSM mass from light curve modelling reported by Aster et al. (2026), Hu et al. (2026) and Shi et al. (2026).
@@ 5.1 SN 2024abvb as a thermonuclear supernova
In Fig. 12, we compare the optical lightcurve of SN 2024abvb to a number of Type Ia templates (Nugent et al., 2002), shifted so that the maxima of their lightcurves coincide. It is clear that a normal Type Ia SN could not be hidden under the light curve of SN 2024abvb, as from around 20 days onward, we would see a difference in the light curve. Moreover, the SN would dominate the emission at the epoch of our second XShooter spectrum, and so we would expect to see underlying photospheric features from the SN. From a luminosity point of view, accommodating a sub-luminous SN 1991bg-like SN is more plausible; in such a scenario, the underlying SN would be significantly fainter than our observed light curve for all epochs where we have spectra, and approximately consistent with our upper limits on the SN magnitude at early and late times. Nevertheless, the presence of H in the CSM would prove challenging to explain. Similarly, a SN Iax could also be hidden below the CSM interaction as they are several magnitudes below both the detections and upper limits of SN 2024abvb at all phases (e.g. Foley et al., 2013; Srivastav et al., 2020) with peak magnitude down to 13-13 mag (SN 2024vjm, Kwok et al., 2025; Zimmerman et al., 2026) or even fainter (SN 2021fcg, Karambelkar et al., 2021). However, they are not found in remote locations and show an association to young stellar populations (Lyman et al., 2018). A similar argument applies to a .Ia SN, whose intrinsic light curve would be hidden beneath the luminosity produced by CSM interaction (e.g. Inserra et al., 2015; Dong et al., 2022; Padilla Gonzalez et al., 2024). Such SNe have been found in remote locations and are thought to be triggered by a detonation of a helium layer on a low-mass CO white dwarf, which would also accommodate the presence of He in the CSM.
@@ 5.3 SN 2024abvb from a massive star or binary system
Figure 13: Proposed common-envelope evolutionary pathway leading to the nested CSM shells observed in SN 2024abvb. Successive mass-loss episodes driven by orbital evolution and outflow precession during the CE/post-CE phase produce concentric toroidal shells with changing symmetry axes, culminating in an electron-capture supernova whose ejecta interact with each shell in sequence.
Figure 14: Schematic of the nested, asymmetric circumstellar medium around SN 2024abvb. Left: the projected geometry of concentric He/C/O toroidal shells at differing orientations, together with the co-located diffuse H-rich envelope. Right: the observed polarimetric evolution, with epoch-coloured markers indicating which shell geometry dominates the scattering at each phase.
@@
Finally, the relative asymmetries we retrieve can be reconciled with a time-dependent mass-loss axis. In a binary/common-envelope context, the orientation and degree of asymmetry of successive ejection episodes can change because of orbital evolution, episodic disk formation and dissipation, or torque-induced precession of an accretion disk or outflow-launching structure (e.g. Ivanova et al., 2013; Tauris et al., 2015). Such mechanisms naturally produce concentric but non-coaxial tori whose apparent flattening and inclination vary with epoch; the outermost shells then record an earlier geometry, while inner shells trace later, possibly reoriented mass-loss. We therefore consider a physically conservative interpretation in which (a) repeated, binary-driven ejection episodes produce concentric toroidal shells, (b) transient colliding-wind shocks enable dust condensation in some episodes (providing the wavelength-dependent polarisation), and (c) changes in the instantaneous mass-loss symmetry axis (e.g. driven by precession or secular orbital reorientation) produce the observed sequence of inclinations and asymmetries. Fig. 13 and 14 show a schematic CE pathway leading to the formation of the SN 2024abvb CSM configuration and a represetation of the nested, asymmetric CSM surrounding SN 2024abvb, respectively. While WR binaries provide a direct observational analogue for nested dusty shells (Lau et al., 2022), detailed hydrodynamic modelling targeted at CE/post-CE mass loss in stripped progenitors is required to quantify whether the exact sequence of asymmetries and kinematics seen here can be reproduced.
@@ 6 Conclusions
We propose the following conservative interpretation. The data are consistent with episodic, binary-driven mass loss from a common envelope scenario (see Fig. 13) leading to the explosion of an electron-capture supernova surrounded by multiple, concentric (ring-like) CSM tori with differing orientations. In this picture (see Fig. 14), the outer (earlier) tori are more oblate, while intermediate shells are more spherical and the innermost shells are again more aspherical, a sequence that can be produced by changes in the instantaneous mass-loss axis driven by orbital evolution, transient accretion disks, or torque-induced precession during common-envelope / post-common-envelope phases. Dust condensed in earlier, more distant tori can naturally explain the late-time, wavelength-dependent increase in polarisation (blue-weighted) when different dusty shells come into the line of sight or are illuminated by the fading transient. This interpretation is supported by empirical examples of nested dust/ring structures in massive colliding-wind systems and the qualitative consistency of our measured kinematics and polarimetric evolution with episodic tilted ejection. While, to our knowledge, this represents the first SN with spectroscopic evidence consistent with nested, concentric tori, detailed hydrodynamic and radiative-transfer models targeted at CE/post-CE mass loss in stripped progenitors are required to confirm whether the precise sequence of inclinations, asymmetries and kinematics can be reproduced.
@@ Appendix A Additional Tables and Figures
Figure 15: Same as Fig.6 but for O ii.
Figure 16: Same as Fig.6 but for C ii.
Figure 17: Same as Fig.6 but for H and including the $\phi=55.7$ spectrum for NIR lines.
@@
Figure 18: The Serkowski fit to the BVRIBVRI ISP estimated based on the FORS2 broadband polarimetry. The best-fitting model (thick solid line), as well as model fits to randomised 1000 Monte Carlo resamplings, are shown.